OpenStudy (anonymous):

Write the equation that satisfies the given conditions. The slope is -4 and the line passes through the point (-2, 3).

OpenStudy (anonymous):

The answer choices are y = -4x - 5 y = -4x + 5 y = -4x + 11

OpenStudy (anonymous):

y = -4x -5

OpenStudy (anonymous):

since the slope (m) is given which is -4 the equation for point slope form is Y=mx+b. You have m, so it becomes y = -4x+b, b is the y-intercept so (x,y) the y is your 3 in (-2,3) and your x is -2. so the equation is now 3 = -4(-2) + b \[3=-8+b \] add 8 to both sides and you get \[11=b\] so you answer \[y=-4x+11\]

OpenStudy (anonymous):

since the slope is -4/1 = -8/2 and you have to cross (-2,3), start from (-2,3) on the graph and go down 8 spaces and 2 spaces to the right, which is going to the y-intercept

OpenStudy (anonymous):

or you could plug in values to check. plug in x=-2 and check whether you get 3 as your y. first one y =3, second y=13, third y=19. so it's the first equation